# Connected Network Model for the Mechanical Loss of Amorphous Materials

### v2

@@ IV.4 Energy Distribution: Barabási-Albert Networks
Figure 5: Average mechanical loss $Q^{-1}\mathcal{V}C/(\beta\gamma_{0}^{2})$ from $1000$ node BA networks (red) and TLS (black). Here energies are drawn from a uniform distribution between 00 and $E_{\rm max}$, with $E_{\rm max}=0.1,0.5,0.9$eV for solid, dished, dotted, and dot-dashed curves respectively. Barriers between states are constrained to be larger than the energy minima with a maximum of $0.2$eV. Results are averaged over 100 random networks.
@@ V Amorphous silicon and titanium dioxide
Several generic properties of the mechanical loss can be inferred directly from Eq. (19). For low frequencies $\omega\tau\ll 1$ the mechanical loss scales as $\omega$, while for high frequencies $\omega\tau\gg 1$ it scales as $\omega^{-1}$. In the intermediate, regime there is a plateau region with peaks corresponding to $\omega\tau_{j}\sim 1$ for the relaxation modes that dissipate significant energy. To calculate mechanical loss, we use a typical value for the elastic modulus $C=50$GPa [18, 19, 20], bare transition rate $k_{0}=10^{13}{\rm s}^{-1}$ [17, 6], and estimate the deformation potential from differences in stress between connected inherent structures [5] as outlined in Supplemental Material III. We emphasize that these assumption are made for calculations within both the TLS and connected network models, and changing the value of $k_{0}$ merely leads to an overall shift in the frequency $\omega$.
@@
The inverse quality factor estimated from the connected network and TLS models found in Fig. 6 have several qualitative and quantitative differences. It can be appreciated that in general, the TLS model underestimates the mechanical loss in comparison to the more comprehensive connected network models. For a-Si within the frequency range relevant for gravitational wave detection, the network model predicts non-monotonic behavior and a peak at $\sim 1$Hz, while the TLS model predicts a peak at $\sim 1,000$ Hz and a mechanical loss that is overall smaller by two orders of magnitude. For a-TiO2, the contrast is even more extreme as the TLS model would suggest vanishing mechanical loss in the same frequency range. This is due to an abundance of very small barriers that concentrate the mechanical loss in the high frequency regime.
@@
It is difficult to compute the mechanical loss (19) at low temperatures since it requires diagonalizing large and nearly singular matrices. However, we can still understand the qualitative temperature dependent behavior by noting that low frequency modes are excited at high temperature (large barrier and large thermal fluctuations) and high frequency modes at low temperature (small barriers and small thermal fluctuations). This implies that at fixed frequency (e.g. $10^{4}$z) the TLS model predicts vanishing mechanical loss above $\sim 300$K for TiO2 while the connected network has several high temperature relaxation modes.
@@ VI Discussion
Motivated by the observation of a connected network of inherent structures coupled by thermally activated transitions in samples of amorphous solids, we have developed a rigorous generalization of the 50 year old TLS model of mechanical loss. Taking a nonequilibrium thermodynamic perspective, the dynamics of the discrete state network of inherent structures is described by a master equation (eq. (8)) for the time dependent probabilities, ultimately leading to an explicit expression for the energy dissipated in a single cycle of an acoustic oscillation (eq. (7)).
The mechanical loss predicted in Fig. 6 at $300$K and $1$kHz are of order $Q_{CN}^{-1}\sim 10^{-1}$ and $\sim 10^{0}$ for a-Si and a-TiO2 respectively. Experimental estimates for the mechanical loss of a-Si can vary significantly depending on film thickness and preparations conditions, ranging from $\sim 10^{-3}$ for poorly annealed to $\sim 10^{-5}$ for well annealed samples [21, 22, 20]. This is a crucial factor to take into account when comparing our numerical predictions quantitatively to experiments, since the timescales accessible in MD simulations are significantly shorter than experimental protocols.
The necessarily fast quench rates in MD simulations result in poorly annealed samples which we expect to increase the overall mechanical loss of our in silico samples. The effect of aging and annealing has been the subject of several recent studies [23, 24, 25, 26]. Indeed, it has been shown that the density of TLS can be reduced by two orders of magnitude when comparing hyperquenched to ultrastable glass models [23]. A promising direction for future study is therefore to explore how the effects of aging and annealing on mechanical loss manifests in the network structure of the materials.
@@
We prepare samples via melt-quench at a quench rate of $10^{11}K/s$ and find inherent structures by thermal search trajectories at $600K$ for $200$ps with a sampling frequency of $100$fs. We determine candidate TLS based on changes in the minimum energy and filter them based on participation ratio and maximum atomic displacement to remove unlikely candidates. Duplicate pairs of TLS are determined and removed based on a root-mean-squared atomic displacement criterion between structures less than $10^{-4}$Å for a-Si. For a-TiO2 we use a displacement criterion of $10^{-3}$Å  and an additional energy difference criterion of $10^{-3}$eV. This procedure is common for amorphous sample preparation [41] and TLS calculations from MD simulation [6, 5].
To determine connected structures, we add an additional step in our analysis and use the same method for removing duplicate TLS to determine which inherent structures are identical. In this way, we connect TLS together (e.g. TLS A-B and B-C becomes A-B-C), thus forming connected networks of inherent structures as shown in Fig. 1. For a-Si we run 100 search trajectories and find networks of $\sim 1000-3000$ inherent structures, and for TiO2 we find considerably fewer and so we ran $200$ search trajectories and find networks with $\sim 10-1000$ inherent structures. We would like to emphasize that our search is by no means an exhaustive search of all the inherent structures. Further simulation details can be found in Supplemental Material III.
Acknowledgements
This research was supported in part by the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. Computational resources and services were provided by Advanced Research Computing at the University of British Columbia and the Digital Research Alliance of Canada (alliancecan.ca).
@@ I Elastic Energy
In this section, we present an alternate derivation of the quality factor (19) based on the phase-lag of the average energy of the system. As discussed in the main text (2), the total energy of the system is the sum of the elastic energy and the average energy of the network
$\mathcal{U}_{\rm tot}=\mathcal{U}_{\rm elastic}+\mathcal{U}_{\rm CN}\ .$ (S1)
Substituting in the linear elastic energy Eq. (2), energy of the connected network Eq. (3), and assuming the approximation of Eq. (13) we have
@@ II TLS derivation
$R=k_{0}\begin{bmatrix}-e^{-\beta(V-E_{1})}&e^{-\beta(V-E_{2})}\\ e^{-\beta(V-E_{1})}&-e^{-\beta(V-E_{2})}\end{bmatrix}\ .$ (S9)
@@
$R=k_{0}e^{-\beta V}\begin{bmatrix}-1&e^{\beta\Delta}\\ 1&-e^{\beta\Delta}\end{bmatrix}\ ,$ (S10)
@@
$M^{-1}=\frac{1}{1+e^{-\beta\Delta}}\begin{bmatrix}1&1\\ -e^{-\beta\Delta}&1\end{bmatrix}\ .$ (S12)
@@ III Simulation details
In this section we provide additional simulation details. By injecting random initial velocities, ten amorphous samples of 1000 atoms each of a-Si and 750 atoms each of TiO2 are prepared by rapidly melting diamond silicon and rutile TiO2 from $100$K to $4000$K in $200$ps, equilibrating at $4000$K for $200$ps, and subsequently cooling to $300$K at a rate of $10^{11}$ K/s in the isobaric ensemble (target pressure P=0). Final amorphous configurations are found by energy minimization of the melt-quenched structure at constant volume.
For a-Si we have an average simulation box length of $27.43\pm 0.01$Å, the density of our samples is $2.261{\rm g/cm^{3}}\pm 0.001$. Using a cut off of $2.9$Å the silicon atoms in the $10$ samples have average coordination (with standard error) $c_{3}=0.42\pm 0.08\%$, $c_{4}=95.8\pm 0.2\%$, $c_{5}=3.8\pm 0.2\%$, and $c_{6}=0.03\pm 0.02\%$, where $c_{i}$ is the percentage of the sample with coordination $i$. This corresponds to $\sim 4\%$ defects in our samples. For a-TiO2 we have a simulation box of lengths $22.97\times 22.97\times 14.79\pm 0.01$Å, for an average density of $4.247{\rm g/cm^{3}}\pm 0.001$.
Once the samples have been prepared, we perform $100$ random thermal searches for a-Si and 200 for a-TiO2 per sample at $600$K for $200$ps. Every $0.1$ps we save the structure of the system and the $2000$ structures per search ($200,000$ per sample a-Si,$400,000$ per sample a-TiO2) are quenched to 00K providing an inherent structure of the system. Sequentially visited structures are considered as candidate connected pairs, and the atomic participation ratio (number of atoms involved in the transition) and maximum atomic displacement between these pairs is calculated and used to filter out unlikely candidates as shown in Fig. S1.
@@
For a-Si, we consider all the candidates with participation ratio $<100$ and $d_{\rm max}>0.1$Å. For a-TiO22{2}2, we do not observe as clear a cut off, so we use a generous cut off of participation ratio $<200$ and $d_{\rm max}>0.01$Å  (Fig. S2). We then remove all duplicate pairs of atomic structures based on total root-mean squared atomic displacement of $10^{-4}$Å  for a-Si. We found that a-TiO2 required a root-mean squared atomic displacement of $10^{-3}$Å  with an additional energy difference criteria of $\Delta E<10^{-3}$eV to identify duplicate structures (Fig. S3).
@@
Figure S3: Distribution of energy difference and total root-mean-squared (RMS) displacement between inherent structures of a-TiO2. The region enclosed by dotted red lines ($|\Delta E|<10^{-3}$eV and RMS displacement $<10^{-3}$Å) are considered the same atomic structure.
@@
An example histogram of the product of longitudinal deformation potentials for one sample of a-Si is shown in Fig. S4 and one sample of TiO2 in Fig. S5. States whose energy deforms in the same direction (increase or decrease energy) will have a positive product, and opposite directions a negative product. Similar to previous studies [6] we observe a fairly wide range of deformation potentials, with the product reaching up to $40{\rm eV}^{2}$ in Fig.  S4.

### v3

@@ IV.4 Energy Distribution: Barabási-Albert Networks
Figure 5: Average mechanical loss $Q^{-1}\mathcal{V}C/(\beta\gamma_{0}^{2})$ from $1000$ node BA networks (red) and TLS (black). Here energies are drawn from a uniform distribution between 0 and $E_{\rm max}$, with $E_{\rm max}=0.1,0.5,0.9$eV for solid, dashed, and dotted curves respectively. Barriers between states are constrained to be larger than the energy minima with a maximum of 0.20.2eV. Results are averaged over 100 random networks.
@@ V Amorphous silicon and titanium dioxide
Several generic properties of the mechanical loss can be inferred directly from Eq. (19). For low frequencies $\omega\tau\ll 1$ the mechanical loss scales as $\omega$, while for high frequencies $\omega\tau\gg 1$ it scales as $\omega^{-1}$. In the intermediate, regime there is a plateau region with peaks corresponding to $\omega\tau_{j}\sim 1$ for the relaxation modes that dissipate significant energy. To calculate mechanical loss, we use a typical value for the elastic modulus C=50C=50GPa [34, 12, 26], bare transition rate $k_{0}=10^{13}{\rm s}^{-1}$ [42, 21], and estimate the deformation potential from differences in stress between connected inherent structures [8] as outlined in Supplemental Material III. We emphasize that these assumptions are made for calculations within both the TLS and connected network models, and changing the value of $k_{0}$ merely leads to an overall shift in the frequency $\omega$.
@@
The inverse quality factor estimated from the connected network and TLS models found in Fig. 6 have several qualitative and quantitative differences. It can be appreciated that in general, the TLS model underestimates the mechanical loss in comparison to the more comprehensive connected network models. For a-Si within the frequency range relevant for gravitational wave detection, the network model predicts non-monotonic behavior and a peak at $\sim 10^{3}$ Hz, while the TLS model predicts a peak at $\sim 1$ Hz and a mechanical loss that is overall smaller by two orders of magnitude. For a-TiO2, the contrast is even more extreme as the TLS model would suggest vanishing mechanical loss in the same frequency range. This is due to an abundance of very small barriers that concentrate the mechanical loss in the high frequency regime.
@@
It is difficult to compute the mechanical loss (19) at low temperatures since it requires diagonalizing large and nearly singular matrices. However, we can still understand the temperature behavior qualitatively by noting that low frequency modes are excited at high temperature (large barrier and large thermal fluctuations) and high frequency modes at low temperature (small barriers and small thermal fluctuations). This implies that at fixed frequency (e.g. $10^{4}$Hz) the TLS model predicts vanishing mechanical loss above $\sim 300$K for TiO2 while the connected network has several high temperature relaxation modes (see Fig. 6).
@@ VI Discussion
Motivated by the observation of a connected network of inherent structures coupled by thermally activated transitions in samples of amorphous solids, we have developed a rigorous generalization of the 50 year old TLS model of mechanical loss. Taking a nonequilibrium thermodynamic perspective, the dynamics of the discrete state network of inherent structures is described by a master equation (eq. (8)) for the time dependent probabilities, ultimately leading to an explicit expression for the energy dissipated in a single cycle of an acoustic oscillation (eq. (7)). The connected network model can describe hypercubic networks that would be formed by entirely independent TLS, but generalizes to TLS interacting via elastic or electrostatic forces and interacting multi-level systems in general.
The mechanical loss predicted in Fig. 6 at 300300K and 11kHz are of order $Q_{CN}^{-1}\sim 10^{-1}$ and $\sim 10^{0}$ for a-Si and a-TiO2 respectively. Experimental estimates for the mechanical loss of a-Si can vary significantly depending on film thickness and preparations conditions, ranging from $\sim 10^{-3}$ for poorly annealed to $\sim 10^{-5}$ for well annealed samples [22, 15, 26]. This is a crucial factor to take into account when comparing our numerical predictions quantitatively to experiments, since the timescales accessible in MD simulations are significantly shorter than experimental protocols.
The necessarily fast quench rates in MD simulations result in poorly annealed samples which we expect to increase the overall mechanical loss of our in silico samples. The effect of aging and annealing has been the subject of several recent studies [17, 23, 16, 33]. Indeed, it has been shown that the density of TLS can be reduced by two orders of magnitude when comparing hyperquenched to ultrastable glass models [17]. A promising direction for future study is therefore to explore how the effects of aging and annealing on mechanical loss manifests in the network structure of the materials. As we have shown, even the smallest networks consisting of three or four states can result in dramatic differences between TLS and connected network predictions, so we expect that even for these well annealed samples it will be important to correctly account for the connectivity between states.
@@
We prepare samples via melt-quench at a quench rate of $10^{11}K/s$ and find inherent structures by thermal search trajectories at $600K$ for 200200ps with a sampling frequency of 100100fs. We determine candidate TLS based on changes in the minimum energy and filter them based on participation ratio and maximum atomic displacement to remove unlikely candidates. Duplicate pairs of TLS are determined and removed based on a root-mean-squared atomic displacement criterion between structures less than $10^{-4}$Å for a-Si. For a-TiO2 we use a displacement criterion of $10^{-3}$Å  and an additional energy difference criterion of $10^{-3}$eV. This procedure is common for amorphous sample preparation [10] and TLS calculations from MD simulation [21, 8].
To determine connected structures, we add an additional step in our analysis and use the same method for removing duplicate TLS to determine which inherent structures are identical. In this way, we connect TLS together (e.g. TLS A-B and B-C becomes A-B-C), thus forming connected networks of inherent structures as shown in Fig. 1. For a-Si we run 100 search trajectories and find networks of $\sim 1000-3000$ inherent structures, and for TiO2 we find considerably fewer and so we ran $200$ search trajectories and find networks with $\sim 10-1000$ inherent structures. Compared to previous studies [8, 21], we report more connected inherent structures for three reasons: i) we allow for sequential transitions that diffuse the system further from the initial structure in state-space, ii) we do not filter based on asymmetries or barriers since all can contribute meaningfully to the mechanical loss in the connected network model, and iii) building on these previous studies, our thermal search trajectories were optimized for finding a large number of connected inherent structures. Further simulation details can be found in Supplemental Material III.
We would like to emphasize that our search is by no means an exhaustive search of all the inherent structures. We have performed convergence tests for the fraction of the network that is connected and the fraction of cycles within the network and find them to be statistically robust (Supplemental Material IV).
Acknowledgments
We thank Jess McIver, Jeff Young, and Ke Zou for many discussions of this work and a critical reading of the manuscript. This research was supported in part by the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program, the New Frontiers in Reserch Fund (Exploration stream), and the Discovery Grant program of the Natural Sciences and Engineering Research Council of Canada. Computational resources and services were provided by Advanced Research Computing at the University of British Columbia and the Digital Research Alliance of Canada (alliancecan.ca).
@@ I Elastic Energy
In this section, we present an alternate derivation of the quality factor (19) based on the phase-lag of the average energy of the system. As discussed in the main text (2), the total energy of the system is the sum of the elastic work and the average energy of the network
$\mathcal{U}_{\rm tot}=\mathcal{W}_{\rm elastic}+\mathcal{U}_{\rm CN}\ .$ (S1)
Substituting in the linear elastic work Eq. (2), energy of the connected network Eq. (3), and assuming the approximation of Eq. (13) we have
@@ II TLS derivation
$R=k_{0}\begin{bmatrix}-e^{-\beta(V-E_{1})}&e^{-\beta(V-E_{2})}\\ e^{-\beta(V-E_{1})}&-e^{-\beta(V-E_{2})}\end{bmatrix}\$. (S9)
@@
$R=k_{0}e^{-\beta V}\begin{bmatrix}-1&e^{\beta\Delta}\\ 1&-e^{\beta\Delta}\end{bmatrix}\$, (S10)
@@
$M^{-1}=\frac{1}{1+e^{-\beta\Delta}}\begin{bmatrix}1&1\\ -e^{-\beta\Delta}&1\end{bmatrix}\$. (S12)
@@
Figure S1: Idealized example of two isolated bond-hopping transitions (left) resulting in four inherent structures A-D. These inherent structures are connected in a symmetric four-state loop, which when driven by $\boldsymbol{\Gamma}=\tfrac{1}{2}(\gamma_{\alpha}\boldsymbol{v}_{\alpha}+\gamma_{\beta}\boldsymbol{v}_{\beta})$ produces identical mechanical loss (scaled dissipation $Q^{-1}\mathcal{V}C/\gamma_{0}^{2}$) with the TLS (black) or connected network (purple) models.
[…]
$v_{0}$ $=[1,1,1,1]\ ,\ \lambda_{0}=0\ ,$ (S13)
$v_{\alpha}$ $=[1,-1,-1,1]\ ,\ \lambda_{\alpha}\ ,$ (S14)
$v_{\beta}$ $=[1,1,-1,-1]\ ,\ \lambda_{\beta}\ ,$ (S15)
$v_{\gamma}$ $=[1,-1,1,-1]\ ,\ \lambda_{\gamma}=\lambda_{\alpha}+\lambda_{\beta}\ .$ (S16)
If the two transitions are independent, then they couple independently to incident mechanical waves and we have
$\Gamma_{A}$ $=\frac{1}{2}\left(\Gamma_{1}+\Gamma_{3}\right),$ (S17)
$\Gamma_{B}$ $=\frac{1}{2}\left(\Gamma_{1}+\Gamma_{4}\right),$ (S18)
$\Gamma_{C}$ $=\frac{1}{2}\left(\Gamma_{2}+\Gamma_{4}\right),$ (S19)
$\Gamma_{D}$ $=\frac{1}{2}\left(\Gamma_{2}+\Gamma_{3}\right)\ .$ (S20)
Since $\boldsymbol{\Gamma}\cdot\boldsymbol{v}_{\gamma}=\Gamma_{A}-\Gamma_{B}+\Gamma_{C}-\Gamma_{D}=0$, the third correlated relaxation mode does not contribute to the mechanical loss and we recover the TLS model description. This coupling is estimated directly from simulations for a-Si and a-TiO2 calculations.
@@ III Simulation details
In this section we provide additional simulation details. By injecting random initial velocities, ten amorphous samples of 1000 atoms each of a-Si and 750 atoms each of TiO2 are prepared by rapidly melting diamond silicon and rutile TiO2 from 100100K to 40004000K in 200200ps, equilibrating at 40004000K for 200200ps, and subsequently cooling to 300300K at a rate of $10^{11}$ K/s in the isobaric ensemble (target pressure P=0). Final amorphous configurations are found by energy minimization of the melt-quenched structure at constant volume.
For a-Si we have an average simulation box length of $27.43\pm 0.01$Å, the density of our samples is $2.261{\rm g/cm^{3}}\pm 0.001$. Using a cut off of 2.92.9Å the silicon atoms in the $10$ samples have average coordination (with standard error) $c_{3}=0.42\pm 0.08\%$, $c_{4}=95.8\pm 0.2\%$, $c_{5}=3.8\pm 0.2\%$, and $c_{6}=0.03\pm 0.02\%$, where $c_{i}$ is the percentage of the sample with coordination $i$. This corresponds to $\sim 4\%$ defects in our samples. For a-TiO2 we have a simulation box of lengths $22.97\times 22.97\times 14.79\pm 0.01$Å, for an average density of $4.247{\rm g/cm^{3}}\pm 0.001$.
Once the samples have been prepared, we perform $100$ random thermal searches for a-Si and 200 for a-TiO2 per sample at 600600K for 200200ps. Every 0.10.1ps we save the structure of the system and the $2000$ structures per search ($200,000$ per sample a-Si,$400,000$ per sample a-TiO2) are quenched to 0K providing an inherent structure of the system. Sequentially visited structures are considered as candidate connected pairs, and the atomic participation ratio (number of atoms involved in the transition) and maximum atomic displacement between these pairs is calculated and used to filter out unlikely candidates as shown in Fig. S2.
@@
For a-Si, we consider all the candidates with participation ratio <100<100 and $d_{\rm max}>0.1$Å. For a-TiO2{2}, we do not observe as clear a cut off, so we use a generous cut off of participation ratio <200<200 and $d_{\rm max}>0.01$Å  (Fig. S3). We then remove all duplicate pairs of atomic structures based on total root-mean squared atomic displacement of $10^{-4}$Å  for a-Si. We found that a-TiO2 required a root-mean squared atomic displacement of $10^{-3}$Å  with an additional energy difference criteria of $\Delta E<10^{-3}$eV to identify duplicate structures (Fig. S4).
@@
Figure S4: Distribution of energy difference and total root-mean-squared (RMS) displacement between inherent structures of a-TiO2. The region enclosed by dotted red lines ($|\Delta E|<10^{-3}$eV and RMS displacement $10^{-3}$Å) are considered the same atomic structure.
@@
An example histogram of the product of longitudinal deformation potentials for one sample of a-Si is shown in Fig. S5 and one sample of TiO2 in Fig. S6. States whose energy deforms in the same direction (increase or decrease energy) will have a positive product, and opposite directions a negative product. Similar to previous studies [21] we observe a fairly wide range of deformation potentials, with the product reaching up to $40{\rm eV}^{2}$ in Fig.  S5.
@@
IV Convergence tests
Finally, although our search for connected networks is by no means exhaustive, we have performed convergence tests for some key statistical properties and find them to be relatively robust. For a-Si (Fig. S7) we find that the fraction of the network that is connected converges to $\sim 0.05\%$ with relatively small spread between our 10 samples. The fraction of four-state cycles in the networks has a larger spread between samples, but remains relatively flat as the fraction of the total network included in the subsample approaches one. For higher order cycles, even faster convergence is observed. Comparing the degree distribution at logarithmically spaced fractions of the total network, we find that as the subsamples approach the full network, it always displays a scale-free structure (power law distribution).
Similar behavior is observed for the a-TiO2 samples (Fig. S8); however, with significantly larger variation between samples and less stable convergence despite doubling the number of trajectories compared to a-Si. This is indicative of the sampling issues we faced. In general, we expect the qualitative trends to be robust to the incomplete sampling of the energy landscape and sampling biases, but quantitative predictions will likely change with improved sampling.
Figure S7: Convergence tests for the statistical properties of our a-Si samples. The fraction of the total network is the number of subsampled nodes divided by the number of nodes in the full network. Each sample is shown as a faded solid curve with the average a completely opaque dashed curve. The percent of the network that is connected is defined as the number of connections divided by the maximum possible number of connections (all-all connected network).
Figure S8: Convergence tests for the statistical properties of a-TiO2 samples, as in Fig. S7.